# Towards Robust Recommendation: A Review and an Adversarial Robustness Evaluation Library

### v1

@@ 8.1.3 Model Module
Model Module serves as the core component of ShillingREC, comprising the Attack Model and Rec Model. Within the Attack Model segment, ShillingREC categorizes shilling attack models into those tailored for explicit feedback data and those for implicit feedback data. Specifically, for explicit feedback data, the shilling attack model generates fake data in the form of (user-item-rating), while for implicit feedback data, it generates fake data in the form of (user-item). In the Rec Model segment, ShillingREC supports various common recommendation models such as MF[92], NeuMF[93], LightGCN[94], and SGL[46], with the ability to easily incorporate additional models. Furthermore, ShillingREC adapts different recommendation models for tasks such as rating prediction and top-k ranking, aiming to achieve a relatively fair evaluation in various experimental settings.
8.1.4 Execution Module
@@ 8.2 Experiment
This section presents an evaluation of the attack performance of several typical adversarial attack models using ShillingREC in scenarios involving rating data attacks or implicit feedback data attacks.
TABLE V: Evaluation results of attacks under the implicit feedback data .
ML-100K ML-1M CiaoDVD
Attack Model Rec Model NDCG@10 NDCG@50 HR@10 HR@50 NDCG@10 NDCG@50 HR@10 HR@50 NDCG@10 NDCG@50 HR@10 HR@50
None MF 0 0.0002 0.0000 0.0011 0 0.0005 0.0003 0.0076 0 0.0004 0 0.0022
NeuMF 0 0 0 0 0.0001 0.0011 0.001 0.0156 0 0.0045 0 0.0216
LightGCN 0 0 0 0 0.0001 0.0011 0.0013 0.0158 0 0 0 0
SGL 0 0 0 0 0.0001 0.0013 0.0012 0.0173 0 0 0 0
Random Attack MF 0.0004 0.0004 0.0011 0.0011 0.0002 0.0007 0.0013 0.0071 0.0075 0.0164 0.0108 0.0519
NeuMF 0.0047 0.0101 0.0090 0.0349 0.0012 0.0037 0.0077 0.0365 0 0.1823 0 0.9892
LightGCN 0.0029 0.0048 0.0045 0.0135 0.0004 0.0014 0.0029 0.0145 0.0077 0.0214 0.0195 0.0801
SGL 0 0 0 0 0.0001 0.0008 0.001 0.0109 0 0.0012 0 0.0065
Bandwagon Attack MF 0.0000 0.0004 0.0000 0.0023 0.0000 0.0002 0.0002 0.002 0.0014 0.0071 0.0022 0.0303
NeuMF 0.0099 0.0195 0.0214 0.0665 0.0010 0.0028 0.0049 0.0264 0.7099 0.7131 0.9632 0.9784
LightGCN 0.0007 0.0022 0.0023 0.0101 0.0002 0.0009 0.0015 0.0108 0.0291 0.0484 0.0606 0.1494
SGL 0 0.0002 0 0.0011 0.0002 0.001 0.0013 0.0128 0 0.0021 0 0.0108
AUSH MF 0.0750 0.1576 0.1781 0.5434 0.0323 0.1354 0.0945 0.4456 0.0142 0.0470 0.0303 0.1861
NeuMF 0.0537 0.1203 0.1206 0.4295 0.0031 0.0123 0.0101 0.0577 0 0.1902 0 0.9892
LightGCN 0.0010 0.1002 0.0034 0.5096 0.0004 0.0017 0.0015 0.0129 0.0449 0.0725 0.0844 0.2100
SGL 0.0045 0.0132 0.0113 0.0541 0.0002 0.0023 0.0008 0.0180 0.0007 0.0016 0.0022 0.0065
RevAdv MF 0.0422 0.1142 0.1116 0.4329 0.0001 0.0009 0.0005 0.0062 0.0155 0.0327 0.0303 0.1126
NeuMF 0.046 0.1072 0.1003 0.3844 0.0008 0.0031 0.0044 0.0272 0.2749 0.2988 0.6320 0.7381
LightGCN 0.0003 0.0592 0.0011 0.3078 0.0003 0.0022 0.0020 0.0242 0.1421 0.2178 0.4004 0.6905
SGL 0.0306 0.0879 0.0643 0.3281 0.0009 0.0041 0.0045 0.0286 0.2798 0.2958 0.3420 0.4134
TABLE VI: Evaluation Results of NeuMF on Explicit Feedback Data.
ML-100K ML-1M CiaoDVD
Attack model MAE RMSE MAE RMSE MAE RMSE
None 0.7539 0.9577 0.6907 0.8845 0.7250 0.9560
Random Attack 0.7564 0.9577 0.6975 0.8915 0.7803 1.0008
Average Attack 0.7581 0.9625 0.6973 0.8919 0.7443 0.9686
LoveHate Attack 0.7546 0.9606 0.6928 0.8859 0.7779 0.9978
AUSH 0.7623 0.9710 0.7058 0.9041 0.7462 0.9785
TABLE VII: Evaluation results for attack on explicit feedback data in AUSH.
ML-100K ML-1M CiaoDVD
REC Model MAE RMSE MAE RMSE MAE RMSE
MF 0.8007 1.0045 0.6936 0.8866 0.7972 1.0889
NeuMF 0.7623 0.9710 0.7058 0.9041 0.7462 0.9785
LightGCN 0.7671 0.9761 0.6914 0.8821 1.0394 1.3632
SGL 0.7851 0.9994 0.7091 0.9112 1.0417 1.3883
8.2.1 Explicit feedback data
To evaluate the capability of the Shilling attack against explicit feedback data, we conducted experiments using four attack models: Random Attack, Average Attack, Love/Hate Attack, and AUSH. We compared the performance differences of the recommender system rating prediction task before and after the attacks. The experimental setup is as follows: we selected three commonly used rating datasets in the recommender system domain, namely ML-100k, ML-1M, and CiaoDVD, for experimentation. For different attack methods, we set the same attack parameters. Specifically, the attacker size was set to 20% of the total number of users in the training dataset; the target item number was set to one third of the attacker size, with popular items selected as targets, and the filler item size was set to the dataset’s average. The victim model used was the classic NeuMF model, and we evaluated the performance of the NeuMF model on the rating prediction task. The specific experimental results are shown in Table VI.
To evaluate the robustness of the recommendation models, we selected AUSH as the attack model and the MF, NeuMF, LightGCN and SGL models as the victim models. We compared the performance of these three models in the rating prediction task under attack scenarios on the ML-100k, ML-1M and CiaoDVD datasets. We utilized MAE and RMSE as evaluation metrics. The specific experimental results are presented in Table VII.
8.2.2 Implicit feedback data
For implicit feedback data, we employed four adversarial attack methods: RandomAttack, BandwagonAttack, AUSH, and RevAdv, to assess the recommender system’s performance on recommending the target item. Evaluation metrics including NDCG and HitRatio were utilized. The experimental setup is as follows: for each attack method, identical attack parameters were set, specifically, the attacker size was set to 20% of the total number of users in the training dataset; the target item number was set to 0.5% of the attacker size, with randomly selected items as targets, and the filler item size was set to the dataset’s average. The victim models employed were MF, NewMF, LightGCN, and SGL, with pairwise optimization strategies utilized to enhance victim model robustness. Experimental results are presented in Table V.
8.2.3 Experimental Analysis
The effectiveness of adversarial attacks on recommender systems is sensitive to the size of the dataset. Generally, adversarial attacks tend to yield better results on small-scale datasets compared to large-scale ones. Compared to supervised recommendation models, SGL, which incorporates unsupervised signals, exhibits superior robustness. This superiority can be attributed to the unsupervised signals filtering out anomalous data during the learning process. In comparison to heuristic methods such as Random attack, attack methods based on optimization and model-based techniques (e.g., AUSH and RevAdv) demonstrate superior efficacy.
9 Future research direction and open issues
@@
From a data perspective, research on adversarial robustness in recommendation algorithms lacks real-world data of fraudulent users. Currently, academic studies are based on artificially simulated fake user data. This setup cannot guarantee the effectiveness of robust recommendation algorithms developed in actual production environments. Furthermore, due to constraints such as privacy concerns in recommender systems and platform rules, shilling attack models in recommender systems suffer from limited attack content. Currently, they can only simulate user-item-rating data and cannot construct other types of simulations. Moreover, treating adversarial attacks as an upstream task in the adversarial robustness of recommender systems may to some extent constrain the research on robust recommendation algorithms.
From a methodological point of view, fraud detection is a crucial approach to improving the adversarial robustness of recommender systems. However, methods based on fraud detection struggle to avoid sample misclassification when aiming at high recall rates. This misclassification can significantly degrade the performance of subsequent training in the recommendation model. Joint training of fraud detection tasks with recommendation tasks presents a viable solution. Works such as GraphRfi [63] and PDR [95] have explored this area and achieved promising results.
While there is a significant body of research on handling natural noise in recommender systems, it is often scenario-specific and does not comprehensively address challenges such as data imbalance, sampling bias, and natural noise. This highlights the need for developing more universally robust recommendation algorithms to tackle these multifaceted issues.
10 Conclusion
@@
In this survey, we provide a comprehensive overview of robustness in recommender systems. We summarize and categorize a large body of existing robustness methods into two main types: adversarial robustness and non-adversarial robustness. In the section on adversarial robustness of recommender systems, we delve into attack and defense perspectives, detailing the classification methods and basic principles of shilling attacks and defenses, along with an introduction to classical methods. In the non-adversarial robustness section, we explore the causes and corresponding methods for non-adversarial robustness issues stemming from data sparsity, natural noise, and data imbalance. Additionally, we present a comprehensive analysis and introduction of commonly used datasets and evaluation metrics in robustness research of recommender systems. To ensure fair and reasonable evaluation of existing shilling attack and defense methods in recommender systems, we propose an adversarial robustness evaluation library–ShillingREC and we experimentally assess several attack and recommendation models within this library. . Finally, we analyze potential issues and future research directions in the field of recommender system robustness.

### v3

@@ VIII-A3 Model Module
Model Module serves as the core component of ShillingREC, comprising the Attack Model and Rec Model. Within the Attack Model segment, ShillingREC categorizes shilling attack models into those tailored for explicit feedback data and those for implicit feedback data. Specifically, for explicit feedback data, the shilling attack model generates fake data in the form of (user-item-rating), while for implicit feedback data, it generates fake data in the form of (user-item). In the Rec Model segment, ShillingREC supports various common recommendation models such as MF[95], NeuMF[96], LightGCN[97], and SGL[49], with the ability to easily incorporate additional models. Furthermore, ShillingREC adapts different recommendation models to two tasks: rating prediction and top-k ranking, aiming to achieve a relatively fair evaluation in various experimental settings.
VIII-A4 Execution Module
@@ VIII-B Experiment
This section presents an evaluation of the attack performance of several typical adversarial attack models using ShillingREC in scenarios involving explicit data attacks or implicit feedback data attacks.
VIII-B1 Explicit Feedback Data
To evaluate the performance and robustness of recommendation models under shilling attacks in explicit feedback scenarios, we conduct rating prediction experiments on four widely used datasets: ML-100K, ML-1M, CiaoDVD, and Epinions. The evaluated recommendation models include MF, NeuMF, LightGCN, and SGL, covering both classic and state-of-the-art methods. The attack models include Random Attack, Average Attack, Love/Hate Attack—widely adopted classical heuristic-based methods—and AUSH, a representative GAN-based attack. All attacks are configured with a consistent setup: the attacker size is set to 20% of the training users; target items are selected from unpopular items, with the number of target items set to 0.5% of the attacker size; and the number of filler items equals the average number of ratings per user in the dataset. We use MAE and RMSE to evaluate recommendation performance, and PS (Prediction Shift) to evaluate robustness. Results are presented in Table V.
VIII-B2 Implicit Feedback Data
To evaluate the performance and robustness of recommendation models under shilling attacks in implicit feedback scenarios, we conduct Top-K recommendation experiments on four widely used datasets: Gowalla, LastFM, ML-1M, and CiaoDVD. Gowalla and LastFM are native implicit feedback datasets, while ML-1M and CiaoDVD are converted from explicit to implicit feedback by applying a rating threshold. The evaluated recommendation models include MF, NeuMF, LightGCN, CDAE, SGL, and SGDL, covering both classic and state-of-the-art methods. The attack models include Random Attack and Bandwagon Attack—widely adopted classical heuristic-based methods—AUSH, a representative GAN-based attack, and RevAdv, a representative optimization-based attack. The attack configurations follow the same setup as in the explicit feedback experiments. We use NDCG and Hit Ratio to evaluate recommendation performance, and T-NDCG and T-HR to evaluate robustness, with all evaluation metrics computed under the Top-K recommendation setting, where K=50. Results are presented in Table VI.
VIII-B3 Experimental Analysis
First, the effectiveness of shilling attacks is sensitive to dataset size. Generally, attacks are more successful on small-scale datasets. For example, in explicit feedback scenarios, attack models achieve higher PS scores on the ML-100K dataset than on the larger ML-1M. This is because user interactions are sparser in smaller datasets, making models more vulnerable to the biased signals from fake users. In contrast, large datasets contain richer and more diverse user behaviors, where real data dominates and weakens the impact of injected fake data.
Second, in implicit feedback scenarios, models using unsupervised signals—such as SGL—show stronger robustness than supervised ones. Under AUSH attacks, SGL consistently outperforms LightGCN across all datasets in robustness metrics (T-NDCG and T-HR). This is due to the filtering effect of unsupervised contrastive learning, which enforces consistency and improves model stability and generalization under shilling attacks.
Lastly, regarding attack methods, optimization-based and model-based attacks (e.g., AUSH and RevAdv) generally outperform heuristic ones like Random Attack and Bandwagon Attack. On the LastFM dataset, for instance, AUSH consistently achieves better attack results across all models. This is because such methods don’t rely on fixed rules but generate more realistic and effective fake behaviors, leading to stronger manipulation.
IX Future Research Direction and Open Issues
@@
From a data perspective, research on adversarial robustness in recommendation algorithms lacks real-world data of fraudulent users. Currently, academic studies are based on artificially simulated fake user data. This setup cannot guarantee the effectiveness of robust recommendation algorithms developed in actual production environments. Furthermore, due to limitations imposed by privacy concerns and platform regulations, the features of fake users generated by the attack model are restricted. Currently, it can only simulate user-item-rating triplet data and cannot replicate the complex features present in real-world scenarios. Moreover, treating adversarial attacks as an upstream task in the adversarial robustness of recommender systems may to some extent constrain the research on robust recommendation algorithms.
From a methodological point of view, there exists a trade-off between the effectiveness and efficiency of adversarial robust attack models. Generally, the more knowledge the attack model possesses about the recommender system, the better the attack’s effectiveness; however, this often results in slower generation of fake data. In current applications, attacks are typically performed as offline tasks with low real-time requirements. Consequently, most existing research has focused primarily on the effectiveness of attack models, with limited attention given to the timeliness of adversarial attacks. In fraud detection models, aiming for a high recall rate can lead to sample misclassification, which may negatively impact subsequent recommendation model training. Joint training of fraud detection and recommendation tasks is a promising approach, as explored in works like GraphRfi [65] and PDR [98]. Similarly, in robust algorithms for adversarial defense strategies, performance and efficiency also present a trade-off. For example, methods based on contrastive learning and unsupervised strategies may introduce additional parameters, reducing the speed of inference. Given that recommendation models require rapid online inference, it is essential to achieve better robustness within acceptable timeliness limits. Therefore, researching high-timeliness robust recommendation algorithms is a critical direction.
From the perspective of model generalization, adversarial robustness largely depends on the design of the attack model, the scale of the attack, and the selection of attack targets. Most existing defense methods are tailored to specific attack strategies or base model designs, and currently, there is no sufficiently robust defense strategy that can protect various types of recommendation models against all forms of attack. Regarding non-adversarial robustness, current research primarily focuses on denoising techniques for natural noise. However, these methods are often limited to narrow scenarios and fail to address issues like data imbalance, sampling bias, and natural noise simultaneously. Thus, developing a general-purpose robust recommendation algorithm is a crucial research focus in recommendation system robustness.
X Conclusion
@@
In this survey, we provide a comprehensive overview of robustness in recommender systems. We summarize and categorize a large body of existing robustness methods into two main types: adversarial robustness and non-adversarial robustness. In the section on adversarial robustness of recommender systems, we delve into attack and defense perspectives, detailing the classification methods and basic principles of shilling attacks and defenses, along with an introduction to classical methods. In the non-adversarial robustness section, we explore the causes and corresponding methods for non-adversarial robustness issues stemming from data sparsity, natural noise, and data imbalance. Additionally, we present a comprehensive analysis and introduction of commonly used datasets and evaluation metrics in robustness research of recommender systems. To ensure fair and reasonable evaluation of existing shilling attack and defense methods in recommender systems, we propose an adversarial robustness evaluation library–ShillingREC and we experimentally assess several attack and recommendation models within this library. Finally, we analyze potential issues and future research directions in the field of recommender system robustness.